# Steiner Traveling Salesman Problem with Time Windows and Pickup–Delivery: integrating classical and quantum optimization

### v2

@@ 1 Introduction
The Vehicle Routing Problem (VRP) is a fundamental combinatorial optimization problem in logistics and transportation. It focuses on determining the most efficient routes for a fleet of vehicles to deliver goods to a set of customers, starting and ending at one or more depots [ghiani2003real, braekers2016vehicle]. Among the most studied extensions of the VRP is the variant involving time windows, where each customer must be served within a specified time interval. This reflects realistic delivery scenarios, where users often select preferred delivery slots when placing an order. These temporal constraints add realism to the model but also significantly increase its computational complexity, as feasible solutions must simultaneously satisfy both spatial and temporal requirements [pugliese2017last, HAM20181, di2020trucks].
Another practically relevant variant is the VRP with pickup and deliveries (VRP-PD), in which each request involves a pair of locations, one for pickup and one for delivery, subject to precedence and capacity constraints. This model reflects the growing complexity of modern logistics systems, where vehicles must not only deliver but also collect items, such as in reverse logistics, product returns or shared mobility services. VRP-PD based models have a wide range of applications, including ride-sharing, e-commerce, hospital logistics and urban freight distribution [li2001metaheuristic, desaulniers2002vrp, ropke2006adaptive].
@@
Gate-based systems perform universal quantum computation by applying sequences of quantum logic gates to qubits, enabling algorithms such as Shor’s factorization [shor1994algorithms] and Grover’s search [grover1996fast], which theoretically provide exponential or quadratic speed-ups over classical methods. In practice, however, these algorithms require large-scale, fault-tolerant architectures with thousands of error-corrected qubits, resources that remain far beyond current technological capabilities [preskill2018quantum]. As a result, gate-based methods are mostly limited to proof-of-concept demonstrations and demand specialized expertise in quantum circuit design, state manipulation, and error correction, restricting their use in applied optimization contexts.
QA, on the other hand, represents a more specialized and technologically mature approach tailored for optimization. It exploits quantum tunneling and adiabatic evolution to search for the minimum of an energy landscape encoded as an Ising or Quadratic Unconstrained Binary Optimization (QUBO) model [glover2018tutorial, glover2020quantum]. This paradigm is particularly appealing because many combinatorial optimization problems can, in principle, be formulated as QUBOs, and modern hardware platforms such as D-Wave’s quantum annealers already allow the execution of optimization instances. QA has therefore been explored in a wide range of combinatorial optimization problems, including scheduling [venturelli2016job, perez2024solving, ciacco2025Educational], routing [holliday2025advanced], facility location [ciacco2026facility, malviya2023logistics], and packing [de2022hybrid, garcia2022comparative].
Despite these efforts, practical limitations persist. The associated energy landscapes often exhibit high ruggedness and numerous local minima, which can hinder the convergence of QA and variational algorithms [abbas2024challenges]. Recent analyses have also emphasized the gap between demonstrating feasibility of quantum formulations and achieving any meaningful notion of quantum advantage. In particular, Smith-Miles et al. [smith2025travelling] argue that current quantum optimization pipelines remain highly sensitive to modeling choices, penalty parameter tuning, hardware noise, and embedding constraints, and therefore exhibit no clear scaling benefits over state-of-the-art classical methods. Their conclusions reinforce the need for a cautious and honest interpretation of existing QA results, especially in routing problems where instance sizes remain small.
@@
The remaining of this paper is structured as follows. In Section 2, we review the scientific literature related to STSP, VRP-TWPD and quantum approaches for combinatorial optimization problems. Section 3 provides a detailed description of the problem under study, including its main characteristics and motivation. Section 6 reports the computational experiments, including the instance generation process, solver setup, and numerical analysis performed with both classical and quantum approaches. Section 7 summarizes the concluding remarks and prospects for future studies.
@@ 2.1.1 Works on the STSP
The STSP was introduced by Cornuéjols et al. [cornuejols1985traveling] and Fleischmann [fleischmann1985cutting]. Cornuéjols et al. [cornuejols1985traveling] presented the Graphical Traveling Salesman Problem, a generalization of the TSP defined on arbitrary graphs, where nodes and edges can be traversed multiple times. The goal is to find a tour that visits all nodes at least once, allowing for repeated traversals. This formulation is particularly suited for structured environments with constrained connectivity, such as city maps or factory layouts. Fleischmann [fleischmann1985cutting] proposed an exact solution method for the STSP based on a cutting plane approach. His model only considers the subset of required nodes and relaxes the requirement of Hamiltonian cycles, allowing for shorter tours. The formulation introduces valid inequalities tailored to the STSP to strengthen the linear relaxation, enabling exact solution of instances on realistic road networks. Letchford et al. [letchford2013compact] introduced three compact formulations for the STSP, leveraging techniques developed for the classical TSP. These include the single-commodity flow, multi-commodity flow, and time-staged formulations. Computational results indicate that the multi-commodity formulation is especially powerful, capable of solving instances with over 200 nodes using standard branch-and-bound algorithms. Notably, it achieves the same linear programming lower bound as Fleischmann’s exponential model, representing a significant computational advancement. Rodríguez-Pereira et al. [rodriguez2019steiner] proposed a recent compact ILP formulation based on the observation that any optimal solution to the STSP traverses each edge at most twice. Their model employs two sets of binary variables to represent the first and second traversal of each edge, significantly reducing model size while preserving accuracy. This formulation enables exact solution of instances with up to 500 nodes, outperforming prior approaches. Interian and Ribeiro [interian2017grasp] developed a metaheuristic based on the Greedy Randomized Adaptive Search Procedure, enhanced with path-relinking and periodic restarts. Their approach incorporates a reduced 2-opt neighborhood, problem-specific greedy criteria, and local search intensification. The use of path-relinking allows structured exploration of elite solution paths, improving overall solution quality. Experimental results confirm its effectiveness on large, sparse STSP instances. Álvarez-Miranda et al. (2019) [alvarez2019note] proposed transforming the STSP into a standard TSP by computing the metric closure over the required nodes, where each edge represents the shortest-path distance in the original graph.
@@ 2.1.2 Works on the VRP-TWPDP
Dumas et al. (1991) [dumas1991pickup] proposed a foundational exact method for the VRP-TWPDP, incorporating vehicle capacity, pickup and delivery precedence, and time window constraints. Their approach is based on a column generation scheme applied to a set-partitioning formulation, where each column represents a feasible route. The subproblem, formulated as a resource-constrained shortest path, is solved via dynamic programming and provides strong lower bounds. Chang et al. (2003) [chang2003real] extended the VRP-TWPDP into a real-time environment, introducing a formulation where customer requests arrive dynamically and must be handled immediately. The authors proposed a heuristic composed of three main components: route construction, route improvement using an anytime algorithm, and a tabu search procedure. Parameters for the tabu search are optimized via Taguchi orthogonal arrays. Their results show that the method is efficient in reducing operational costs under real-time constraints. Both approaches provide solid foundations for practical vehicle routing under temporal and resource constraints. Fabri and Recht (2006) [fabri2006dynamic] proposed a dynamic extension of the VRP-TWPDP in which each customer request includes two separate time windows, one for pickup and one for delivery. Departing from the stretch-factor approach in [caramia2002routing], their model allows explicit time window constraints and vehicle waiting. The solution combines exact single-vehicle routing via dynamic programming and a fast heuristic to allocate requests across vehicles.
@@
The problem considered in this work, however, presents a substantially richer structure. The addition of time windows combined with pickup-and-delivery operations introduces temporal feasibility constraints and precedence relationships that significantly alter the underlying combinatorial landscape. A related line of research is the VRP-TWPD, which represents the closest classical counterpart. Early exact methods, such as those by Dumas et al. [dumas1991pickup], already demonstrated how the interaction between time-window feasibility and pickup–delivery pairing severely restricts tractable instance sizes, with exact algorithms typically limited to a few dozen requests. Subsequent heuristic and metaheuristic studies achieve larger scales, up to a few hundred requests, yet consistently highlight the intrinsic difficulty of jointly managing precedence, time windows, and routing decisions.
@@
Cornuéjols et al. [cornuejols1985traveling] theoretical
Fleischmann [fleischmann1985cutting] 292 nodes exact
Letchford et al. [letchford2013compact] 200 nodes exact
Rodríguez-Pereira et al. [rodriguez2019steiner] 500 nodes exact
Interian and Ribeiro [interian2017grasp] 3353 nodes heuristic
Álvarez-Miranda et al. [alvarez2019note] 300 nodes exact
Dumas et al. [dumas1991pickup] 15 nodes exact
Chang et al. [chang2003real] 100 nodes heuristic
Fabri and Recht [fabri2006dynamic] 1000 nodes heuristic
@@ 2.2 Quantum Approaches
A broad range of combinatorial optimization problems have been explored through QA. For a comprehensive and structured discussion of existing QA applications, we refer the reader to the surveys by Ciacco et al. [ciacco2025review], Osaba et al. [osaba2022systematic], Yulianti and Surendro [yulianti2022annealing] and Muhamediyeva et al. [muhamediyeva2026optimization]. Ciacco et al. [ciacco2025review] offer a detailed examination of quantum algorithm design across multiple domains, including healthcare, finance, production planning, and logistics, together with a systematic classification of methodological frameworks. Osaba et al. [osaba2022systematic] present a unified and end-to-end review of eighteen years of research (2004–2021) at the intersection between quantum computing and routing problems, analysing 53 contributions and identifying emerging trends and open challenges. In addition, Yulianti and Surendro [yulianti2022annealing] provide a systematic review of QA implementations, mapping 229 studies across thirteen application areas and four methodological families. Their work emphasizes how QA research has expanded in parallel with the evolution of commercial quantum annealers, highlighting both performance gains and persistent hardware-driven limitations. Muhamediyeva et al. [muhamediyeva2026optimization] examines the practical limitations of both quantum annealing and gate-based frameworks, and provides an accessible discussion of the quantum approximation optimization and quantum alternating operator ansatz paradigms.
Within this wider landscape, the only quantum-based study explicitly addressing the STSP is presented by Ciacco et al. [ciacco2025steiner]. In their work, the authors reformulate the STSP as a QUBO model derived from an integer linear programming formulation, introducing a dedicated preprocessing phase aimed at removing redundant arcs and reducing the dimensionality of the problem before its quantum optimisation. The resulting model is evaluated using both the D-Wave LeapBQM simulator and a Quantum Processing Unit (QPU), demonstrating that the preprocessing step substantially enhances computational performance and contributes to obtaining higher-quality solutions.
@@
Tambunan et al. [tambunan2022quantum] investigate QA approaches to the VRP with Weighted Road Segments, formulating the problem as a QUBO and solving it on a D-Wave QPU. Le et al. [le2023quantum] address the Orienteering Problem, another routing-related variant, through a QUBO formulation executed on D-Wave systems. Sinno et al. [sinno2023performance] provide an extensive empirical evaluation of commercial QA platforms on the Capacitated VRP (CVRP), conducting over 30 hours of experiments using D-Wave hybrid solvers. Their findings reveal that while QPU time remains extremely small, solution quality degrades with increasing constraint density rather than with problem size alone. In particular, the authors highlighting how modelling choices, and especially the number and structure of constraints in the QUBO or CQM formulation, critically influence performance. Further contributions include the work of Mori and Furukawa [10.3389/fphy.2023.1129594], who introduce the Adjuster Routing Problem, a disaster-response VRP in which insurance companies must deploy large numbers of adjusters across affected areas. Their QUBO formulation incorporates multiple operational constraints, and numerical experiments on D-Wave demonstrate that both solution quality and computational effort are highly sensitive to parameter selection. Recent contributions have further extended QA to realistic logistics settings. Osaba et al. [osaba2024solving] consider advanced delivery settings involving heterogeneous vehicle fleets, delivery priorities, and dual capacity constraints (weight and volume). Their experiments on real-world instances using D-Wave’s hybrid CQM solver confirm the practical relevance of quantum–classical methods for operational logistics. Mario et al. [mario2024hybrid] propose two hybrid classical-quantum strategies for real-time route optimisation. Their methods decompose the CVRP into clustering and routing phases, using classical Fuzzy C-Means for capacitated clustering and QA for solving either CVRP or TSP subproblems. Osaba et al. [osaba2025package] investigate the VRP-TWPD, integrating realistic constraints such as simultaneous pickup–delivery operations, time-window restrictions, and vehicle-specific mobility limitations. Their results, obtained through D-Wave’s Hybrid CQM Leap framework on seven classes of instances, reinforce the feasibility and robustness of hybrid QA approaches for complex logistics and routing environments.
@@
Ciacco et al. [ciacco2025steiner] STSP 9 nodes Hybrid BQM D-Wave Leap
@@
Tambunan et al. [tambunan2022quantum] VRP with Weighted Road Segments 5 vehicles QPU D-Wave
Le et al. [le2023quantum] Orienteering Problem 7 cities QPU D-Wave
Sinno et al. [sinno2023performance] CVRP 80 customers Hybrid CQM D-Wave Leap
Mori and Furukawa [10.3389/fphy.2023.1129594] Adjuster Routing Problem 20 adjusters Hybrid CQM D-Wave Leap
Osaba et al. [osaba2024solving] VRP 29 customers Hybrid CQM D-Wave Leap
Mario et al. [mario2024hybrid] CVRP 37 customers QPU D-Wave
Osaba et al. [osaba2025package] VRP-TWPD 24 nodes Hybrid CQM D-Wave Leap

### v3

@@ 1 Introduction
The Vehicle Routing Problem (VRP) is a fundamental combinatorial optimization problem in logistics and transportation. It focuses on determining the most efficient routes for a fleet of vehicles to deliver goods to a set of customers, starting and ending at one or more depots [22, 4]. Among the most studied extensions of the VRP is the variant involving time windows, where each customer must be served within a specified time interval. This reflects realistic delivery scenarios, where users often select preferred delivery slots when placing an order. These temporal constraints add realism to the model but also significantly increase its computational complexity, as feasible solutions must simultaneously satisfy both spatial and temporal requirements [16, 17].
Another practically relevant variant is the VRP with pickup and deliveries (VRP-PD), in which each request involves a pair of locations, one for pickup and one for delivery, subject to precedence and capacity constraints. This model reflects the growing complexity of modern logistics systems, where vehicles must not only deliver but also collect items, such as in reverse logistics, product returns or shared mobility services. VRP-PD based models have a wide range of applications, including ride-sharing, e-commerce, hospital logistics and urban freight distribution [30, 15, 43].
@@
Gate-based systems perform universal quantum computation by applying sequences of quantum logic gates to qubits, enabling algorithms such as Shor’s factorization [44] and Grover’s search [25], which theoretically provide exponential or quadratic speed-ups over classical methods. In practice, however, these algorithms require large-scale, fault-tolerant architectures with thousands of error-corrected qubits, resources that remain far beyond current technological capabilities [41]. As a result, gate-based methods are mostly limited to proof-of-concept demonstrations and demand specialized expertise in quantum circuit design, state manipulation, and error correction, restricting their use in applied optimization contexts.
QA, on the other hand, represents a more specialized and technologically mature approach tailored for optimization. It exploits quantum tunneling and adiabatic evolution to search for the minimum of an energy landscape encoded as an Ising or Quadratic Unconstrained Binary Optimization (QUBO) model [23, 24]. This paradigm is particularly appealing because many combinatorial optimization problems can, in principle, be formulated as QUBOs, and modern hardware platforms such as D-Wave’s quantum annealers already allow the execution of optimization instances. QA has therefore been explored in a wide range of combinatorial optimization problems, including scheduling [48, 40, 8], routing [26], facility location [11, 31], and packing [14, 21].
Despite these efforts, practical limitations persist. The associated energy landscapes often exhibit high ruggedness and numerous local minima, which can hinder the convergence of QA and variational algorithms [1]. Recent analyses have also emphasized the gap between demonstrating feasibility of quantum formulations and achieving any meaningful notion of quantum advantage. In particular, Smith-Miles et al. [46] argue that current quantum optimization pipelines remain highly sensitive to modeling choices, penalty parameter tuning, hardware noise, and embedding constraints, and therefore exhibit no clear scaling benefits over state-of-the-art classical methods. Their conclusions reinforce the need for a cautious and honest interpretation of existing QA results, especially in routing problems where instance sizes remain small.
@@
The remaining of this paper is structured as follows. In Section 2, we review the scientific literature related to STSP, VRP-TWPD and quantum approaches for combinatorial optimization problems. Section 3 provides a detailed description of the problem under study, including its main characteristics and motivation. Section 4 presents the formal mathematical formulations of the problem, for both the ABF and NBF variants, along with a discussion on their equivalence, dimensionality, and linearization mechanisms. Two simplified problem variants, namely STSP-TW and STSP-PD, are also derived from the general STSP-TWPD formulation. Section 5 introduces the arcs filtering and graph reduction method aimed at improving solution efficiency. Section 6.3 analyzes the dimensionality and computational complexity of the formulations, both before and after the application of the preprocessing method. Section 6 reports the computational experiments, including the instance generation process, solver setup, and numerical analysis performed with both classical and quantum approaches. Section 7 summarizes the concluding remarks and prospects for future studies.
@@ 2.1.1 Works on the STSP
The STSP was introduced by Cornuéjols et al. [12] and Fleischmann [20]. Cornuéjols et al. [12] presented the Graphical Traveling Salesman Problem, a generalization of the TSP defined on arbitrary graphs, where nodes and edges can be traversed multiple times. The goal is to find a tour that visits all nodes at least once, allowing for repeated traversals. This formulation is particularly suited for structured environments with constrained connectivity, such as city maps or factory layouts. Fleischmann [20] proposed an exact solution method for the STSP based on a cutting plane approach. His model only considers the subset of required nodes and relaxes the requirement of Hamiltonian cycles, allowing for shorter tours. The formulation introduces valid inequalities tailored to the STSP to strengthen the linear relaxation, enabling exact solution of instances on realistic road networks. Letchford et al. [29] introduced three compact formulations for the STSP, leveraging techniques developed for the classical TSP. These include the single-commodity flow, multi-commodity flow, and time-staged formulations. Computational results indicate that the multi-commodity formulation is especially powerful, capable of solving instances with over 200 nodes using standard branch-and-bound algorithms. Notably, it achieves the same linear programming lower bound as Fleischmann’s exponential model, representing a significant computational advancement. Rodríguez-Pereira et al. [42] proposed a recent compact ILP formulation based on the observation that any optimal solution to the STSP traverses each edge at most twice. Their model employs two sets of binary variables to represent the first and second traversal of each edge, significantly reducing model size while preserving accuracy. This formulation enables exact solution of instances with up to 500 nodes, outperforming prior approaches. Interian and Ribeiro [27] developed a metaheuristic based on the Greedy Randomized Adaptive Search Procedure, enhanced with path-relinking and periodic restarts. Their approach incorporates a reduced 2-opt neighborhood, problem-specific greedy criteria, and local search intensification. The use of path-relinking allows structured exploration of elite solution paths, improving overall solution quality. Experimental results confirm its effectiveness on large, sparse STSP instances. Álvarez-Miranda et al. (2019) [2] proposed transforming the STSP into a standard TSP by computing the metric closure over the required nodes, where each edge represents the shortest-path distance in the original graph.
@@ 2.1.2 Works on the VRP-TWPDP
Dumas et al. (1991) [18] proposed a foundational exact method for the VRP-TWPDP, incorporating vehicle capacity, pickup and delivery precedence, and time window constraints. Their approach is based on a column generation scheme applied to a set-partitioning formulation, where each column represents a feasible route. The subproblem, formulated as a resource-constrained shortest path, is solved via dynamic programming and provides strong lower bounds. Chang et al. (2003) [6] extended the VRP-TWPDP into a real-time environment, introducing a formulation where customer requests arrive dynamically and must be handled immediately. The authors proposed a heuristic composed of three main components: route construction, route improvement using an anytime algorithm, and a tabu search procedure. Parameters for the tabu search are optimized via Taguchi orthogonal arrays. Their results show that the method is efficient in reducing operational costs under real-time constraints. Both approaches provide solid foundations for practical vehicle routing under temporal and resource constraints. Fabri and Recht (2006) [19] proposed a dynamic extension of the VRP-TWPDP in which each customer request includes two separate time windows, one for pickup and one for delivery. Departing from the stretch-factor approach in [5], their model allows explicit time window constraints and vehicle waiting. The solution combines exact single-vehicle routing via dynamic programming and a fast heuristic to allocate requests across vehicles.
@@
The problem considered in this work, however, presents a substantially richer structure. The addition of time windows combined with pickup-and-delivery operations introduces temporal feasibility constraints and precedence relationships that significantly alter the underlying combinatorial landscape. A related line of research is the VRP-TWPD, which represents the closest classical counterpart. Early exact methods, such as those by Dumas et al. [18], already demonstrated how the interaction between time-window feasibility and pickup–delivery pairing severely restricts tractable instance sizes, with exact algorithms typically limited to a few dozen requests. Subsequent heuristic and metaheuristic studies achieve larger scales, up to a few hundred requests, yet consistently highlight the intrinsic difficulty of jointly managing precedence, time windows, and routing decisions.
@@
Cornuéjols et al. [12] theoretical
Fleischmann [20] 292 nodes exact
Letchford et al. [29] 200 nodes exact
Rodríguez-Pereira et al. [42] 500 nodes exact
Interian and Ribeiro [27] 3353 nodes heuristic
Álvarez-Miranda et al. [2] 300 nodes exact
Dumas et al. [18] 15 nodes exact
Chang et al. [6] 100 nodes heuristic
Fabri and Recht [19] 1000 nodes heuristic
@@ 2.2 Quantum Approaches
A broad range of combinatorial optimization problems have been explored through QA. For a comprehensive and structured discussion of existing QA applications, we refer the reader to the surveys by Ciacco et al. [7], Osaba et al. [39], Yulianti and Surendro [50] and Muhamediyeva et al. [35]. Ciacco et al. [7] offer a detailed examination of quantum algorithm design across multiple domains, including healthcare, finance, production planning, and logistics, together with a systematic classification of methodological frameworks. Osaba et al. [39] present a unified and end-to-end review of eighteen years of research (2004–2021) at the intersection between quantum computing and routing problems, analysing 53 contributions and identifying emerging trends and open challenges. In addition, Yulianti and Surendro [50] provide a systematic review of QA implementations, mapping 229 studies across thirteen application areas and four methodological families. Their work emphasizes how QA research has expanded in parallel with the evolution of commercial quantum annealers, highlighting both performance gains and persistent hardware-driven limitations. Muhamediyeva et al. [35] examines the practical limitations of both quantum annealing and gate-based frameworks, and provides an accessible discussion of the quantum approximation optimization and quantum alternating operator ansatz paradigms.
Within this wider landscape, the only quantum-based study explicitly addressing the STSP is presented by Ciacco et al. [9]. In their work, the authors reformulate the STSP as a QUBO model derived from an integer linear programming formulation, introducing a dedicated preprocessing phase aimed at removing redundant arcs and reducing the dimensionality of the problem before its quantum optimisation. The resulting model is evaluated using both the D-Wave LeapBQM simulator and a Quantum Processing Unit (QPU), demonstrating that the preprocessing step substantially enhances computational performance and contributes to obtaining higher-quality solutions.
@@
Tambunan et al. [47] investigate QA approaches to the VRP with Weighted Road Segments, formulating the problem as a QUBO and solving it on a D-Wave QPU. Le et al. [28] address the Orienteering Problem, another routing-related variant, through a QUBO formulation executed on D-Wave systems. Sinno et al. [45] provide an extensive empirical evaluation of commercial QA platforms on the Capacitated VRP (CVRP), conducting over 30 hours of experiments using D-Wave hybrid solvers. Their findings reveal that while QPU time remains extremely small, solution quality degrades with increasing constraint density rather than with problem size alone. In particular, the authors highlighting how modelling choices, and especially the number and structure of constraints in the QUBO or CQM formulation, critically influence performance. Further contributions include the work of Mori and Furukawa [34], who introduce the Adjuster Routing Problem, a disaster-response VRP in which insurance companies must deploy large numbers of adjusters across affected areas. Their QUBO formulation incorporates multiple operational constraints, and numerical experiments on D-Wave demonstrate that both solution quality and computational effort are highly sensitive to parameter selection. Recent contributions have further extended QA to realistic logistics settings. Osaba et al. [37] consider advanced delivery settings involving heterogeneous vehicle fleets, delivery priorities, and dual capacity constraints (weight and volume). Their experiments on real-world instances using D-Wave’s hybrid CQM solver confirm the practical relevance of quantum–classical methods for operational logistics. Mario et al. [32] propose two hybrid classical-quantum strategies for real-time route optimisation. Their methods decompose the CVRP into clustering and routing phases, using classical Fuzzy C-Means for capacitated clustering and QA for solving either CVRP or TSP subproblems. Osaba et al. [38] investigate the VRP-TWPD, integrating realistic constraints such as simultaneous pickup–delivery operations, time-window restrictions, and vehicle-specific mobility limitations. Their results, obtained through D-Wave’s Hybrid CQM Leap framework on seven classes of instances, reinforce the feasibility and robustness of hybrid QA approaches for complex logistics and routing environments.
@@
Ciacco et al. [9] STSP 9 nodes Hybrid BQM D-Wave Leap
@@
Tambunan et al. [47] VRP with Weighted Road Segments 5 vehicles QPU D-Wave
Le et al. [28] Orienteering Problem 7 cities QPU D-Wave
Sinno et al. [45] CVRP 80 customers Hybrid CQM D-Wave Leap
Mori and Furukawa [34] Adjuster Routing Problem 20 adjusters Hybrid CQM D-Wave Leap
Osaba et al. [37] VRP 29 customers Hybrid CQM D-Wave Leap
Mario et al. [32] CVRP 37 customers QPU D-Wave
Osaba et al. [38] VRP-TWPD 24 nodes Hybrid CQM D-Wave Leap